a) Find the adjusted reliability of the complete test whose split-half reliability value was found
to be 0.57. What method will be used for this?

b) For the test with a reliability coefficient of 0.87 what will be the coefficient of split-half
reliability for this test?
1
Expert's answer
2020-10-08T10:10:14-0400

a)We can use Kuder-Richardson formula 20 to calculate split-half reliability.

"r_{KR20}=(\frac{k}{k-1})(1-\frac{\sum pq}{\sigma^2}"


b) Formula for calculating Reliability coefficient;


"RC=(\frac{N}{N-1})\times\frac{ (Total Variance-Sum of Variance)}{Total Variance}"